OpenStudy (anonymous):

Area of circle whose circumference is 34pi?

OpenStudy (solomonzelman):

area of the circle is given by the following formula: (where \({\rm A}\) denotes the area, and \(\rm r \) denotes the radius) \( {\rm A} = \pi {\rm r}^2\)

OpenStudy (solomonzelman):

the circumference of the circle is given by the formula: \({\rm C}=2\pi{\rm r}\) (where \(\rm C\) denotes the circumference, and \( {\rm r}\) denotes the radius)

OpenStudy (solomonzelman):

so, you can do this: \({\rm C}=2\pi {\rm r}\) \(\displaystyle \frac{\rm C}{2\pi}= {\rm r}\) (solving for r in terms of C) and then plug it into the area formula. \(\displaystyle \rm A=\pi {\rm r}^2\) \(\displaystyle \rm A=\pi \left( \frac{\rm C}{2\pi} \right)^2\) \(\displaystyle \rm A=\pi \cdot \frac{\rm C^2}{4(\pi^2)} \) \(\displaystyle \rm A=\frac{\rm C^2}{4\pi} \)

OpenStudy (solomonzelman):

So we have shown that Area in term of the circumference is: \(\displaystyle \rm A=\frac{\rm C^2}{4\pi} \)

OpenStudy (solomonzelman):

now, you can plug your circumference that you are given. \(\displaystyle \rm A=\frac{\rm (34\pi)^2}{4\pi} \)

OpenStudy (solomonzelman):

continue from here.

OpenStudy (anonymous):

Thank you! :)

OpenStudy (solomonzelman):

yw